OpenStudy (anonymous):

-1/3x - 1/4y = 1/2 Put the equation in slope-intercept form

OpenStudy (anonymous):

What is the given number for x and y?

OpenStudy (anonymous):

That's all that shows in the question :c But I think: x = -1/3 y = -1/4 Is this what you mean?

OpenStudy (anonymous):

add 1/3x to the left side, then divide by -1/4 or multiply by -4 to get rid of the number on y. You should end up with y=-4/3x-2

OpenStudy (anonymous):

-4/3x-2